carmelle:

A city planner is rerouting traffic in order to work on a stretch of road. The equation of the path of the old route can be described as y = 2/5x − 4. What should the equation of the new route be if it is to be perpendicular to the old route and will go through point (P, Q)? A. y − Q = -5/2(x − P) B. y − Q = 2/5(x − P) C. y − P = -5/2x − Q) D. y − P = 2/5(x − Q)

Vocaloid:

point slope form: y - y1 = m(x-x1) starting with the original line: y = (2/5)x - 4 if the new route is perpendicular, the new slope m must be the negative reciprocal of the old slope 2/5. calculate the new m value. after that, since it passes through (P,Q), you can let these be your (x1,y1) values in the point slope formula